# Attenuation of the ultra-high-energy neutrino flux by dark matter scatterings

### v1

@@ Abstract
A flux of ultra-high-energy (UHE) neutrinos is generally expected to be produced by astrophysical sources at cosmological distances and to reach Earth. In this paper, we investigate the impact of neutrino scattering with dark matter (DM) particles in both the intergalactic medium and the Milky Way on the total flux, energy spectrum, and arrival directions of UHE neutrinos. We emphasize the complementarity of neutrino detectors at different latitudes to probe the anisotropy in the flux at Earth due to the attenuation of the neutrino flux in the Milky Way dark matter halo. We also discuss that, with mild astrophysical assumptions, limits on the DM-$\nu$ scattering cross section can be placed even if the neutrino sources are unknown. Finally, we explore all this phenomenology with the recent UHE neutrino event KM3230213A, and place the corresponding limits in the DM-$\nu$ scattering cross-section.
@@ 1 Introduction
The origin of the ultra-high-energy cosmic rays (UHECRs) remains a mystery to this day. The most accepted explanation is that UHECRs are accelerated in extreme astrophysical environments via electromagnetic processes like Fermi acceleration (for reviews, see e.g. [1, 2]). If this is the case, one expects that the UHECRs could interact with the radiation and the matter surrounding the sources, together with background radiation en-route to Earth, producing a flux of ultra-high energy (UHE) neutrinos that could be detected at Earth.
@@
In the Standard Model of Particle Physics neutrinos interact very weakly with matter and radiation, and the UHE neutrino flux is expected to arrive to Earth isotropically and practically without attenuation. On the other hand, it is now firmly established that the Universe is permeated by an additional component of matter, called Dark Matter (DM). While the interactions of neutrinos with ordinary matter are well described by electroweak theory (at least up to the energies that have been probed), the interactions of neutrinos with dark matter are completely unknown. Upper limits on the dark matter-neutrino cross-section have been derived for $E_{\nu}\sim 100$ eV from the Lyman-alpha forest [3] or from the number of Milky Way dwarf spheroidal galaxies [4]; for $E_{\nu}\sim 10$ MeV from supernova neutrinos [5, 6]; for $E_{\nu}\sim 1$ TeV from the Seyfert Galaxy NGC 1068 [7] and for $E_{\nu}\sim 100$ TeV from the blazar TXS 0506+056 [8, 9]. However, the scattering cross-section for $E_{\nu}\gtrsim 100$ TeV remains totally unconstrainted, save for extrapolations of the existing constraints to larger energies for a postulated energy dependence of the cross-section.
@@
In this paper, we investigate the attenuation of the ultrahigh-energy neutrino flux due to scatterings with dark matter in both the intergalactic medium and the Milky Way halo. Our analysis is conservative since we do not account for potential attenuation effects near the sources of UHE neutrinos. We argue that dark matter–neutrino interactions can lead to significant and distinctive modifications to the observed flux, energy spectrum, and arrival directions of astrophysical neutrinos. As a case study, we focus on the recently reported event KM3-230213A [10], and discuss its consistency with the null results from IceCube and the Pierre Auger Observatory [11, 12].
@@
The paper is organized as follows. In Section 2 we analyze the attenuation of a neutrino flux generated at high redshift due to scatterings with dark matter in the intergalactic medium and in the Milky Way. In Section 3 we apply this formalism to investigate the implications of the attenuation on a flux that saturates the Waxman-Bahcall bound, and on a concrete scenario in which the event KM3-230213A could be part of a diffuse flux. In Section 4 we emphasize the importance of a network of neutrino detectors at different latitudes to establish whether the UHE neutrino flux is affected by attenuation due to dark matter particles. Finally, in Section 5, we present our conclusions.
@@ 2 Neutrino flux attenuation by dark matter scatterings
Let us assume that neutrinos are injected into the Universe at time $t$ with energy $E$, with rate ${\cal I}(E,t)$. The time evolution of the comoving number density of neutrinos is governed by the following transport equation [13]
$\frac{\partial n_{\nu}(E,t)}{\partial t}=\frac{\partial}{\partial E}\left[H(t)En_{\nu}(t,E)\right]+\mathcal{I}(E,t)-n_{\mathrm{DM}}(t)\sigma(E)n_{\nu}(E,t)\,$, (1)
where we have included the neutrino energy redshift due to the expansion of the Universe, with rate $H(t)$; and the decrease in the neutrino number density due to interactions with dark matter particles with density $n_{\rm DM}(t)$ and energy-dependent cross-section $\sigma(E)$. Here we neglect for simplicity the production of a secondary neutrino flux in the scattering process.
The transport equation can be rewritten in terms of the redshift as
@@
which has as solution  [^1] To solve the equation it is convenient to perform the change of variables $\mathcal{Z}(E,z)\equiv(1+z)n\big{(}E(1+z),z\big{)}$ [14].
$n_{\nu}(E,z)=\frac{1}{1+z}\int_{z}^{z_{\rm max}}\frac{\mathrm{d}z^{\prime}}{H(z^{\prime})}\mathcal{I}\left(\frac{1+z^{\prime}}{1+z}E,z^{\prime}\right)\exp\Big{[}-\displaystyle\int_{z}^{z^{\prime}}\frac{\mathrm{d}z^{\prime\prime}}{H(z^{\prime\prime})(1+z^{\prime\prime})}n_{\mathrm{DM}}(z^{\prime\prime})\sigma\left(\frac{1+z^{\prime\prime}}{1+z}E\right)\Big{]}\,$, (3)
where we have assumed that neutrinos are injected only at redshifts $z\leq z_{\rm max}$. The differential neutrino flux at a terrestrial detector (located at $z=0$) can be straightforwardly calculated using that neutrinos move relativistically, so that $d\phi/dE=n_{\nu}(E,0)$:
$\frac{\mathrm{d}\phi}{\mathrm{d}E}(E,z=0)$ $=\int_{0}^{z_{\rm max}}\frac{\mathrm{d}z^{\prime}}{H(z^{\prime})}\mathcal{I}\big{(}(1+z^{\prime})E,z^{\prime}\big{)}\exp\Big{[}-\displaystyle\int_{0}^{z^{\prime}}\frac{\mathrm{d}z^{\prime\prime}}{H(z^{\prime\prime})(1+z^{\prime\prime})}n_{\mathrm{DM}}(z^{\prime\prime})\sigma\big{(}(1+z^{\prime\prime})E\big{)}\Big{]}$
@@
Here, we have introduced for convenience an energy dependent optical depth
@@
where we have expressed the dark matter number density in terms of the mass density $n_{\rm DM}(z)=\rho_{\rm DM}(z)/M_{\rm DM}$ and we have used that the mass density and the expansion rate scale with the redshift as $\rho_{\rm DM}(z)=\rho_{{\rm DM},0}(1+z)^{3}$ and $H(z)=H_{0}\sqrt{\Omega_{\Lambda,0}+\Omega_{\rm m,0}(1+z)^{3}}$. Here, $H_{0}$, $\Omega_{\Lambda,0}$, $\Omega_{{\rm m},0}$ and $\rho_{{\rm DM},0}$ are respectively the Hubble constant, the density parameters of cosmological constant and matter, and the mass density of dark matter, all determined at the current cosmic time. In our numerical analysis we will use $H_{0}=67.7\,\mathrm{km/s/Mpc}$, $\Omega_{\Lambda,0}=0.689$ and $\Omega_{{\rm m},0}=0.311$, $\rho_{{\rm DM},0}=1.26\times 10^{-6}\,\mathrm{GeV/cm^{3}}$ [15].
In the case in which the cross-section has a simple power-law dependence with the energy, $\sigma_{n}(E)=\sigma(E_{0})\left(\frac{E}{E_{0}}\right)^{n}$, one obtains
@@
can be interpreted as an effective column density traversed by a neutrino produced at redshift $z$ when the cross section is proportional to $E^{n}$.
Once the neutrinos enter the Milky Way, their flux is subject to an additional attenuation due to scatterings with DM particles in the halo of our Galaxy. Since the Earth is not located at the center of the Milky Way, the attenuation depends on the arrival direction. Using that the flux at the halo boundary (which we identify with the virial radius $R_{\rm vir}\simeq 197.5$ kpc) is isotropic and given by Eq. (4), the flux at Earth can be calculated from
@@
where the optical depth due to scattering with the DM in the Milky Way can be cast in an analogous way to Eq. (6):
$\tau_{\mathrm{MW}}(\Omega,E)=\frac{\sigma(E)}{M_{\mathrm{DM}}}\ell(E,\alpha,\beta)\;.$ (9)
Here $\ell(E,\alpha,\delta)$ is the dark matter column density traversed by the neutrino flux in the Milky Way in the direction defined by the right ascension $\alpha$ and declination $\delta$ as seen from the Earth:
$\ell(E,\alpha,\delta)\equiv\int_{0}^{R_{\rm h.b.}}\rho_{\mathrm{DM}}\big{(}{\vec{r}}(s,\alpha,\delta)\big{)}\,\mathrm{d}s\,$, (10)
with $\rho_{\mathrm{DM}}(\vec{r})$ the Milky Way DM mass density, $R_{\rm h.b.}$ the halo boundary, and $\vec{r}(s)$ is expressed in Cartesian Galactic coordinates as $\vec{r}(s)=s(\cos\delta\cos\alpha,\cos\delta\sin\alpha,\sin\delta)$ with $s$ the distance along the line of sight. In our analysis we will adopt two possible choices of the dark matter density profile in the Milky Way, corresponding to a cuspy profile and to a cored profile. For the cuspy profile we will adopt the Navarro-Frenk-White profile  [16, 17], given by
$\rho(r)=\frac{\rho_{0}}{\left(\dfrac{r}{r_{s}}\right)\left(1+\dfrac{r}{r_{s}}\right)^{2}}\;,$ (11)
with density parameter $\rho_{0}=1.15\times 10^{7}\,M_{\odot}/\mathrm{kpc}$ and scale radius $r_{s}=14.8\,\mathrm{kpc}$[18] while for the cored profile we will adopt the Burkert profile [19], given by:
$\rho(r)=\frac{\rho_{0}}{\left(1+\dfrac{r}{r_{c}}\right)\left[1+\left(\dfrac{r}{r_{c}}\right)^{2}\right]}\;,$ (12)
@@
We show in Fig. 1 the effective column density $\ell_{n}(z)$ traversed by neutrinos produced at a redshift $z$ for different choices of the energy dependence of the scattering cross-section, n=4,2,0,2n=4,2,0,-2. For comparison, we also show in yellow the ranges of the column density in the Milky Way for different neutrino arrival directions. As apparent from the plot, the effective column density in the intergalactic medium is subdominant (albeit non negligible) if the neutrinos are produced at redshift $z\lesssim 0.2$, while it can be dominant if the neutrinos are produced at redshift $z\gtrsim 1$, especially when the scattering cross-section has a strong dependence with the energy. The effective column density is in the range $10^{22}-10^{25}\,{\rm GeV}/{\rm cm}^{2}$. Therefore, for dark matter masses $M_{\rm DM}=1$ GeV, the attenuation can be sizable for scattering cross-sections larger than the Thomson cross-section $\sigma_{\rm T}=6.6\times 10^{-25}\,{\rm cm}^{2}$. A detailed discussion will be presented in Section 3.
@@
Figure 1: Effective integrated DM column density of a high-redshift source, for different energy-dependent cross sections. The yellow bands indicate the ranges of the DM column density in the Milky Way for a NFW profile or a Burkert profile, depending on the neutrino arrival direction.
@@
The column density in the Milky Way DM halo has a an ${\cal O}(1)$ dependence with the arrival direction, which is shown in Fig. 2 for the NFW profile (left plot) and the Burkert profile (right plot). The column density is largest in the direction of the Galactic Center (with equatorial coordinates $\alpha\simeq 17^{h}45^{m}\simeq 266^{\circ}$ and $\delta\simeq-29^{\circ}$), and smallest in the direction of the Galactic Anticenter. The anisotropy in the column density is stronger for a cuspy profile than for a cored profile, although a cored profile has on average a larger column density. We also include for reference the location of the highest energy neutrino observed to date, KM3-230213A, with arrival direction centered in $\alpha\simeq 94.3^{\circ}$ and $\delta\simeq-7.8^{\circ}$, and a 99% containment region of $3^{\circ}$. radius [21].
@@
Figure 2: Integrated DM column density from the Milky Way at different directions, for an NFW (left) and Burkert (right) profile. The red dot corresponds to the arrival direction of KM3-230213A.
@@
If the neutrino flux is isotropic at the halo boundary, the attenuation due to DM scattering would generate an anisotropic flux at the detector, cf. Eq. (8). In Fig. 3 we show the flux attenuation as a function of the declination, integrating over a day (or equivalently integrating over the right-ascension), and for different values of the cross-section per DM mass. For small values of $\sigma/M_{\rm DM}$, the attenuation due to the Milky Way is negligible and the flux at the Earth’s surface is roughly identical to the flux at the halo boundary. On the other hand, as the cross-section increases, there is a reduction of the flux from the North Celestial Pole ($\delta=90^{\circ}$), and an even larger reduction of the flux from the South Celestial Pole ($\delta=-90^{\circ}$). Moreover, the anisotropy in the flux between these two directions is larger and larger as the cross-section increases. The actual number of events at a detector, on the other hand, depends critically on its location, which could be used to probe the possibility that the UHE neutrino flux could be attenuated. This will be discussed in detail in Section 4.
@@
Figure 3: Neutrino-flux attenuation due to dark matter-neutrino scatterings in the Milky Way halo as a function of the declination, for a NFW profile (solid line) or a Burkert profile (dashed line), averaged over one day (or equivalently averaged over right ascension).

### v2

@@ Abstract
A flux of ultra-high-energy (UHE) neutrinos, produced by astrophysical sources at cosmological distances, is anticipated to exist and reach Earth. In this paper, we investigate the impact on the total flux, energy spectrum, and arrival directions of UHE neutrinos of neutrino-dark matter (DM) scatterings. We study scatterings both in the intergalactic medium and in the Milky Way. We emphasize the complementarity among neutrino detectors at different latitudes, that can probe anisotropies induced by neutrinos scattering with the Milky Way DM halo. We also discuss that, with mild astrophysical assumptions, limits on the DM-$\nu$ scattering cross section can be placed even if the neutrino sources are unknown. Finally, we explore all this phenomenology with the recent UHE neutrino event KM3230213A, and place the corresponding limits on the DM-$\nu$ scattering cross section.
@@ 1 Introduction
The origin of the ultra-high-energy cosmic rays (UHECRs) remains a mystery to this day. The most accepted explanation is that UHECRs are accelerated in extreme astrophysical environments via electromagnetic processes like Fermi acceleration (for reviews, see, e.g., Refs. [1, 2]). If this is the case, UHECRs are expected to interact with the radiation and matter surrounding the sources, as well as with background radiation en-route to Earth, producing a flux of ultra-high energy (UHE) neutrinos that could be detected at Earth.
@@
In the Standard Model of Particle Physics neutrinos interact very weakly with matter and radiation, and the UHE neutrino flux is expected to arrive at Earth isotropically and with negligible attenuation. However, it is now firmly established that the Universe is permeated by an additional component of matter, called Dark Matter (DM). While the interactions of neutrinos with ordinary matter are well-described by electroweak theory (at least up to the energies that have been probed), the interactions of neutrinos with DM are completely unknown. Upper limits on the DM-neutrino cross section have been derived for neutrino energies $E_{\nu}\sim 100\,\mathrm{eV}$} from the Lyman-alpha forest [3] and from the number of Milky Way dwarf spheroidal galaxies [4, 5]; for $E_{\nu}\sim 10$ MeV from supernova neutrinos [6, 7]; for $E_{\nu}\sim 1$ TeV from the Seyfert Galaxy NGC 1068 [8] and for $E_{\nu}\sim 100$ TeV from the blazar TXS 0506+056 [9, 10]. However, the scattering cross section for $E_{\nu}\gtrsim 100$ TeV remains totally unconstrainted, save for extrapolations of other constraints to higher energies assuming a specific energy dependence of the cross section.
@@
In this paper, we investigate the attenuation of the UHE neutrino flux due to scatterings with DM in both the intergalactic medium and the Milky Way halo. Our analysis is conservative, as we do not include potential further attenuation effects near the (so far unknown) sources of UHE neutrinos. We argue that DM–neutrino interactions can lead to significant and unique modifications to the observed flux, energy spectrum, and arrival directions of UHE neutrinos. As a case study, we focus on the recently reported event KM3-230213A [11], including its consistency with null results from IceCube and the Pierre Auger Observatory [12, 13]. We show that, with mild astrophysical assumptions, this observation sets limits on DM-neutrino scattering even if its origin is unknown.
@@
The paper is organized as follows. In Section 2 we analyze the attenuation of a neutrino flux generated at high redshift due to scatterings with DM in the intergalactic medium and in the Milky Way. In Section 3 we apply this formalism to investigate the implications of the attenuation of a diffuse UHE flux, arguing that the Waxman-Bahcall bound allows to set limits on neutrino-DM scattering even if the unattenuated flux is unknown. We apply this idea to a concrete scenario in which the event KM3-230213A could be part of a diffuse flux, setting the corresponding limit. In Section 4 we emphasize the importance of a network of neutrino detectors at different latitudes to robustly establish whether the UHE neutrino flux is affected by attenuation due to DM particles. Finally, in Section 5, we present our conclusions.
@@ 2 Neutrino flux attenuation by dark matter scatterings
Let us assume that neutrinos are injected into the Universe at cosmological time $t$ with energy $E$, with comoving rate ${\cal I}(E,t)$. The time evolution of the comoving number density of neutrinos is governed by the following transport equation [14]
$\frac{\partial n_{\nu}(E,t)}{\partial t}=\mathcal{I}(E,t)+\frac{\partial}{\partial E}\left[H(t)En_{\nu}(t,E)\right]-n_{\mathrm{DM}}(t)\sigma(E)n_{\nu}(E,t)\,$, (1)
where we have included the neutrino energy redshift due to the expansion of the Universe, with rate $H(t)$; and the decrease in the neutrino number density due to interactions with DM particles with density $n_{\rm DM}(t)$ and energy-dependent cross section $\sigma(E)$. Here we neglect for simplicity the production of a secondary neutrino flux at lower energies in the scattering process. Since the astrophysical neutrino flux is expected to fall with energy, this secondary flux is expected to be small with respect to the more abundant lower-energy primary flux.
The transport equation can be rewritten in terms of redshift as
@@
which has as solution
$n_{\nu}(E,z)=\frac{1}{1+z}\int_{z}^{z_{\rm max}}\frac{\mathrm{d}z^{\prime}}{H(z^{\prime})}\mathcal{I}\left(\frac{1+z^{\prime}}{1+z}E,z^{\prime}\right)\exp\left[-\displaystyle\int_{z}^{z^{\prime}}\frac{\mathrm{d}z^{\prime\prime}}{H(z^{\prime\prime})(1+z^{\prime\prime})}n_{\mathrm{DM}}(z^{\prime\prime})\sigma\left(\frac{1+z^{\prime\prime}}{1+z}E\right)\right]\,$, (3)
where we have assumed that neutrinos are injected only at redshifts $z\leq z_{\rm max}$. This solution can be readily obtained by the change of variables $\mathcal{Z}(E,z)\equiv(1+z)n\big(E(1+z),z\big)$ [15]. The prefactor $(1+z^{\prime})/(1+z)$ multiplying the neutrino energy $E$ corresponds to energy redshift due to the expansion of the Universe, and the exponential term captures neutrino attenuation due to scattering with DM particles.
The differential neutrino flux at a terrestrial detector (located at $z=0$) can be straightforwardly calculated using that neutrinos move relativistically, so that $d\phi/dE=n_{\nu}(E,0)$:
$\frac{\mathrm{d}\phi}{\mathrm{d}E}(E,z=0)$ $=\int_{0}^{z_{\rm max}}\frac{\mathrm{d}z^{\prime}}{H(z^{\prime})}\mathcal{I}\big((1+z^{\prime})E,z^{\prime}\big)\exp\left[-\displaystyle\int_{0}^{z^{\prime}}\frac{\mathrm{d}z^{\prime\prime}}{H(z^{\prime\prime})(1+z^{\prime\prime})}n_{\mathrm{DM}}(z^{\prime\prime})\sigma\big((1+z^{\prime\prime})E\big)\right]$
@@
Here, we have introduced for convenience an energy-dependent optical depth
@@
where we have expressed the DM number density in terms of the mass density $n_{\rm DM}(z)=\rho_{\rm DM}(z)/M_{\rm DM}$, and we have used that the mass density and the expansion rate scale with redshift as $\rho_{\rm DM}(z)=\rho_{{\rm DM},0}(1+z)^{3}$ and $H(z)=H_{0}\sqrt{\Omega_{\Lambda,0}+\Omega_{\rm m,0}(1+z)^{3}}$. $H_{0}$, $\Omega_{\Lambda,0}$, $\Omega_{{\rm m},0}$ and $\rho_{{\rm DM},0}$ are, respectively, the Hubble constant, the density parameters of cosmological constant and matter, and the mass density of DM; all at the present cosmic time. In our numerical analysis, we use $H_{0}=67.7\,\mathrm{km/s/Mpc}$, $\Omega_{\Lambda,0}=0.689$, $\Omega_{{\rm m},0}=0.311$, and $\rho_{{\rm DM},0}=1.26\times 10^{-6}\,\mathrm{GeV/cm^{3}}$ [16].
If the neutrino-DM cross section has a simple power-law dependence with neutrino energy, $\sigma_{n}(E)=\sigma(E_{0})\left(\frac{E}{E_{0}}\right)^{n}$, one obtains
@@
can be interpreted as an effective column density traversed by a neutrino produced at redshift $z$ when the cross section is proportional to $E^{n}$. The factor $(1+z^{\prime})^{n}$ captures the fact that neutrino energy, and hence the neutrino-DM cross section, depends on redshift.
Once neutrinos enter the Milky Way, their flux is subject to an additional attenuation due to scatterings with DM particles in the halo of our galaxy. Since the Earth is not located at the center of the Milky Way, the attenuation depends on the arrival direction. Using that the flux at the halo boundary (which we identify with the virial radius $R_{\rm vir}\simeq 198$ kpc [17]) is isotropic and given by Eq. (4), the flux at Earth can be calculated from
@@
where the optical depth due to scattering with the DM in the Milky Way can be cast in an analogous way to Eq. (6),
$\tau_{\mathrm{MW}}(\Omega,E)=\frac{\sigma(E)}{M_{\mathrm{DM}}}\ell(\alpha,\beta)\;.$ (9)
Here $\ell(\alpha,\delta)$ is the DM column density traversed by the neutrino flux in the Milky Way in the direction with right ascension $\alpha$ and declination $\delta$ as seen from Earth,
$\ell(E,\alpha,\delta)\equiv\int_{0}^{R_{\rm vir}}\rho_{\mathrm{DM}}\big({\vec{r}}(s,\alpha,\delta)\big)\,\mathrm{d}s\,$, (10)
where $\rho_{\mathrm{DM}}(\vec{r})$ is the Milky Way DM mass density and $\vec{r}(s)$ is expressed in Cartesian Galactic coordinates as $\vec{r}(s,\alpha,\delta)=s\,(\cos\delta\cos\alpha,\cos\delta\sin\alpha,\sin\delta)+\vec{r}_{\odot}$. $s$ is the distance along the line of sight, and $\vec{r}_{\odot}=-r_{\odot}\,(\cos\delta_{\mathrm{GC}}\cos\alpha_{\mathrm{GC}},\cos\delta_{\mathrm{GC}}\sin\alpha_{\mathrm{GC}},\sin\delta_{\mathrm{GC}})$ is the Earth position, with $r_{\odot}=8.18\,\mathrm{kpc}$, $\delta_{\mathrm{GC}}=-28.92^{\circ}$, and $\alpha_{\mathrm{GC}}=266.42^{\circ}$ the coordinates of the Galactic Center as seen from Earth [18]. In our analysis we adopt two possible choices for the DM density profile in the Milky Way, corresponding to a cuspy profile and to a cored profile. For the cuspy profile, we adopt the Navarro-Frenk-White (NFW) profile [19, 20],
$\rho(\vec{r})=\frac{\rho_{0}}{\left(\dfrac{r}{r_{s}}\right)\left(1+\dfrac{r}{r_{s}}\right)^{2}}\;,$ (11)
with density parameter $\rho_{0}=1.15\times 10^{7}\,M_{\odot}/\mathrm{kpc}$ and scale radius $r_{s}=14.8\,\mathrm{kpc}$ [21]; while for the cored profile we adopt the Burkert profile [22],
$\rho(\vec{r})=\frac{\rho_{0}}{\left(1+\dfrac{r}{r_{c}}\right)\left[1+\left(\dfrac{r}{r_{c}}\right)^{2}\right]}\;,$ (12)
@@
We show in Fig. 1 the effective column density $\ell_{n}(z)$ traversed by neutrinos produced at redshift $z$ for energy dependences of the scattering cross section, n=4,2,0,2n=4,2,0,-2. For comparison, we also show in yellow the column density traversed in the Milky Way, where the different values in the shaded area correspond to different neutrino arrival directions. As apparent from the plot, the effective column density in the intergalactic medium is subdominant (albeit non-negligible) if neutrinos are produced at redshift $z\lesssim 0.2$, while it can be dominant if neutrinos are produced at redshift $z\gtrsim 1$, especially when the scattering cross section has a strong dependence with energy. The effective column density is in the range $10^{22}-10^{25}\,{\rm GeV}/{\rm cm}^{2}$. Therefore, for DM masses $M_{\rm DM}=1$ GeV, the attenuation can be sizable for scattering cross sections larger than the Thomson cross section $\sigma_{\rm T}=6.6\times 10^{-25}\,{\rm cm}^{2}$. We present a detailed discussion of phenomenologically relevant cross sections in Section 3.
@@
Figure 1: Effective integrated DM column density of a high-redshift source, for different energy-dependent cross sections. The yellow bands indicate the ranges of the DM column density in the Milky Way for an NFW profile or a Burkert profile, depending on the neutrino arrival direction.
@@
The column density in the Milky Way DM halo depends on the arrival direction, which is shown in Fig. 2 for the NFW profile (left plot) and the Burkert profile (right plot). The column density is largest in the direction of the Galactic Center (with equatorial coordinates $\alpha\simeq 17^{h}45^{m}\simeq 266^{\circ}$ and $\delta\simeq-29^{\circ}$), and smallest in the opposite direction. The anisotropy in the column density is stronger for a cuspy profile than for a cored profile, although a cored profile has on average a larger column density because both profiles have the same total DM mass. We also include for reference the location of the highest-energy neutrino observed to date, KM3‑230213A, with arrival direction centered in $\alpha\simeq 94.3^{\circ}$ and $\delta\simeq-7.8^{\circ}$, and a 99% containment region of $3^{\circ}$ radius [23].
@@
Figure 2: Integrated Milky Way DM column density at different directions, for an NFW (left) and Burkert (right) profile. The red dot corresponds to the arrival direction of KM3‑230213A.
@@
If the neutrino flux is isotropic at the halo boundary, the attenuation due to DM scattering would generate an anisotropic flux at the detector, cf. Eq. (8). In Fig. 3 we show the flux attenuation as a function of the declination, integrating over a day (or, equivalently, integrating over the right ascension), for different values of the cross section per DM mass $\sigma/M_{\mathrm{DM}}$. For small values of $\sigma/M_{\rm DM}$, the attenuation due to the Milky Way is negligible and the flux at Earth is roughly identical to the flux at the halo boundary. In turn, as the cross section increases, there is a reduction of the flux from the North Celestial Pole ($\delta=90^{\circ}$), and an even larger reduction of the flux from the South Celestial Pole ($\delta=-90^{\circ}$). Moreover, the anisotropy in the flux between these two directions becomes larger as the cross section increases. Therefore, the number of events at a given detector that is more sensitive to a part of the sky depends critically on its location. This feature, that has been exploited within a single detector using high-energy astrophysical neutrino data [24, 25], can be used to robustly probe whether the UHE neutrino flux is attenuated by the Milky Way DM. We discuss this in detail in Section 4.
@@
Figure 3: Neutrino-flux attenuation due to DM-neutrino scatterings in the Milky Way halo as a function of the declination, for an NFW profile (solid line) or a Burkert profile (dashed line), averaged over one day (or, equivalently, averaged over right ascension).